I need to find the volume of a cylinder with a base area of 25 inches squared and height equal to the radius and I did that and got the answer. But it says I have to write the answer in terms of inches squared. What's that?
Answer by cleomenius(959)

Volume = A * h
Volume = pi^r^2 *h
2.82 is the radius approx.
So, 25 inches*2 * 2.82 = 70.5 inches cubed.
Volume is a three dimensional measurement, it will be a cubed value. length * Width * height.
Area is two dimensional measurement, it will be a squared value. Lenght * Width.
Cleomenius.